OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). y = −1/3x − 6 y = −1/3x − 14 y = 3x − 6 y = 3x − 14

OpenStudy (anonymous):

okay so you got y by itself, thats a good start, now to have two parallel lines, the slopes of the lines must be the same.

OpenStudy (anonymous):

then you use the point slope form of the equation of a line

OpenStudy (anonymous):

can you work it out for me dont get it at all

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

3x-y=d now just substitute (4,-2) at x any y to get d ...see what you get?

OpenStudy (anonymous):

ohh!

OpenStudy (anonymous):

the answer is d right

OpenStudy (ghazi):

correct

OpenStudy (anonymous):

i have more is that coool

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6). y =1/3 x + 7 y = −3x + 3 y =1/3 x + 3 y = −3x + 7

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

b

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line 2x − 3y = 3 and passes through the point (−8, 2). y = x2/3 − 10 y = x2/3 + 10 y = −3/2x − 10 y = −3/2x + 10

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

d

OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). y = 2x + 6 y = −1/2x + 3 y = −1/2x + 6 y = 2x + 3 @ghazi

OpenStudy (ghazi):

c

OpenStudy (anonymous):

Select the inequality that corresponds to the given graph. 5x + 7 2 and 10x 30 6x + 7 25 or 2x - 8 < -4 4x + 10 6 or 5x - 6 14 37 8x - 3 61

OpenStudy (anonymous):

nvm forget tha one

OpenStudy (anonymous):

-7y - 17 > 11 Part 1: Solve the inequality above. Part 2: Describe the graph of the solution. @ghazi

OpenStudy (ghazi):

y>4

OpenStudy (anonymous):

would you graph to the right with and open circle ?

OpenStudy (ghazi):

 but dude i am not damn sure about this one :)